# Deep Modularity Networks with Diversity–Preserving Regularization

### v1

@@ Abstract
Graph clustering plays a crucial role in graph representation learning but often faces challenges in achieving feature-space diversity. While Deep Modularity Networks (DMoN) leverage modularity maximization and collapse regularization to ensure structural separation, they do not explicitly encourage diversity in the feature space among clusters. We address this limitation by proposing Deep Modularity Networks with Diversity-Preserving Regularization (DMoN–DPR), which introduces three novel regularization terms: distance-based for inter-cluster separation, variance-based for intra-cluster diversity, and entropy-based for balanced assignments. Our method enhances clustering performance on benchmark datasets, namely Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics, achieving significant improvements in Normalized Mutual Information (NMI), and F1 scores. These results demonstrate the effectiveness of incorporating diversity-preserving regularizations in creating meaningful and interpretable clusters, especially in feature-rich datasets.
@@ 1 Introduction
Graph clustering is a crucial problem within the graph representation learning domain, and is essential for various applications, including but not limited to community detection in social networks (Perozzi et al., 2014; Xiao et al., 2015), data exploration (Perozzi and Akoglu, 2018), and functional module identification in biological networks (Jin et al., 2021). Consequently, there has been a surge in methods aimed at enhancing graph clustering performance. Recent advancements have included techniques like Graph Neural Networks (GNNs) (Scarselli et al., 2008), which leverage node features and graph structure for representation learning, often through unsupervised training (Tsitsulin et al., 2023). In this context, graph pooling methods have also become prominent, as they provide a way to coarsen graphs by aggregating nodes into clusters (Cangea et al., 2018). Yet, some of these methods such as DiffPool (Ying et al., 2018) and MinCutPool (Bianchi et al., 2019) were found to be computationally costly, or too rigid, leading to poor convergence (Tsitsulin et al., 2023).
To address these limitations, Deep Modularity Networks (DMoN) (Tsitsulin et al., 2023) were introduced, combining spectral modularity maximization (Newman, 2006) with collapse regularization (Tsitsulin et al., 2023) to avoid trivial clustering solutions. While DMoN improves structural separation, it does not explicitly enforce diversity among clusters in the feature space. This limitation can result in structurally distinct but feature-wise homogeneous clusters, reducing effectiveness in applications requiring diverse and meaningfully differentiated clusters.
@@
To overcome the limitations of DMoN and enhance feature-space diversity in graph clustering, we propose an extension called Deep Modularity Networks with Diversity-Preserving Regularization (DMoN–DPR). Our method augments the original DMoN optimization objective by introducing additional regularization terms specifically designed to promote diversity among clusters. These include a distance-based regularization term, which penalizes clusters with centroids too close in feature space to encourage distinct separation; a variance-based regularization term, which maximizes the variance within each cluster’s assignment probabilities to capture a broader range of node features; and an entropy-based regularization term, which ensures a more balanced and uniform node distribution by maximizing the entropy of cluster assignments.
Our proposed DMoN-DPR method addresses the shortcomings of existing approaches by explicitly enforcing feature-space diversity and ensuring that clusters remain well-separated, all while preventing them from becoming overly homogeneous in their node features. As a result, the clusters produced are both structurally cohesive and feature-wise distinct. Following the evaluation framework in (Tsitsulin et al., 2023), we demonstrate through extensive experiments that our method improves clustering performance on the benchmark datasets, noticeably on Coauthor CS and Coauthor Physics datasets, which benefit from this enriched representation due to their rich feature spaces.
@@ 2 Related Work
Early methods for clustering on graphs often treated node features independently from graph structure. A straightforward example is k-means(features) (Lloyd, 1982), which clusters nodes solely based on attributes and disregards their connectivity. While simple, this approach often overlooks crucial structural relationships. To incorporate at least some graph information, k-means(DeepWalk) (Perozzi et al., 2014) combines node attributes with random-walk-based embeddings. This introduces a basic integration of graph structure but still relies on a two-step process of embedding and clustering. A more advanced alternative is k-means(DGI) (Tsitsulin et al., 2023), which employs node representations learned via Deep Graph Infomax (Veličković et al., 2018) and thereby captures both local and global context. However, because these k-means strategies only merge attributes and structure post hoc—rather than jointly optimizing them—they may lag behind end-to-end trainable methods that unify both graph topology and node features in a single framework.
Moving toward more integrated approaches, Defferrard et al. (2016) introduced one of the first spectral graph convolution methods based on Chebyshev polynomials. Their technique approximates the spectral filters necessary for convolutional operations on graphs, reducing the computational cost of graph Fourier transforms and providing a scalable foundation for subsequent GNN designs. This seminal work illustrated how learned transformations on graph-structured data could drastically improve learning efficiency and accuracy, motivating the community to delve deeper into neural approaches.
Beyond simply learning node-level representations, pooling mechanisms aim to capture hierarchical structures within graphs. Neural Overlapping Community Detection (NOCD) (Shchur and Günnemann, 2019) applies a probabilistic model that directly optimizes the likelihood of reconstructing the graph. Although effective in certain scenarios, NOCD may face scalability challenges and underutilizes node features in complex datasets. One of the earliest neural pooling frameworks, DiffPool (Ying et al., 2018), introduced a soft-clustering strategy that learns cluster assignments in an end-to-end manner. Despite its strong performance, its quadratic time complexity renders it less practical for large-scale graphs (Tsitsulin et al., 2023).
To address computational bottlenecks and overcome potential overfitting, MinCut Pooling (Bianchi et al., 2020) reimagines the clustering problem through a spectral lens, incorporating normalized minCUT objectives with orthogonality constraints. While these constraints can improve cluster quality, they frequently lead to convergence issues. A related simplified variant, Ortho (Bianchi et al., 2020), focuses exclusively on orthogonality regularization, thus discarding many graph-specific objectives. This omission can produce suboptimal results, as the method forgoes some of the structural cues integral to good cluster formation. In another line of work, SAGPool (Lee et al., 2019) introduces a self-attention mechanism that ranks node importance, thereby selecting the most informative nodes during pooling. Although SAGPool effectively combines feature and topology considerations, computing attention coefficients can increase training costs, particularly for large graphs.
Orthogonal to pooling-based approaches, Deep Graph Infomax (DGI) (Velickovic et al., 2019) pioneered an approach that maximizes mutual information between local substructures and a global summary of the graph. This method facilitates robust, label-agnostic node embeddings. The framework was later extended by InfoGraph (Sun et al., 2019), which applies a similar principle at the graph-level, yielding state-of-the-art results for unsupervised and semi-supervised tasks by capturing interactions between subgraphs and a holistic representation.
Most recently, Deep Modularity Networks (DMoN) (Tsitsulin et al., 2023) have demonstrated a powerful alternative by directly incorporating modularity optimization with collapse regularization in an end-to-end trainable pipeline. DMoN bridges traditional graph community detection and neural feature extraction, achieving strong empirical performance on both synthetic and real-world graphs. By jointly optimizing for high modularity and balancing cluster assignment, DMoN outperforms many existing approaches in terms of normalized mutual information (NMI) and modularity scores, thus setting a new standard for clustering-oriented GNNs.
3 Deep Modularity Networks with Diversity–Preserving Regularization
@@
$L_{\text{DMoN}}(C;A)=\underbrace{-\frac{1}{2m}\text{Tr}(C^{\top}BC)}_{\text{Modularity term}}+\underbrace{\sqrt{\frac{k}{n}}\left\|\sum_{i}C_{i}^{\top}\right\|_{F}-1}_{\text{Collapse Regularization term}}$ (2)
where $\frac{1}{2m}\text{Tr}(C^{\top}BC)$ is the modularity term (measuring how well clusters are internally connected compared to random chance), $B=A-\frac{dd^{\top}}{2m}$ is the modularity matrix, $d$ is the degree vector, $m$ is the total number of edges, $\sqrt{\frac{k}{n}}\left\|\sum_{i}C_{i}^{\top}\right\|_{F}-1$ represents the collapse regularization term, which discourages the formation of trivial clusters by penalizing the Frobenius norm of the assignment matrix $C$, $\|\cdot\|_{F}$ denotes the Frobenius norm, and $k$ is the number of clusters. This regularization term encourages balanced cluster assignments, thereby improving the quality of clustering and avoiding degenerate solutions.
Building upon the DMoN framework, we augment the DMoN objective function $L_{\text{DMoN}}$ with three new regularization terms, resulting in the following enhanced objective function:
$L_{\text{DMoN--DPR}}(C;A)=L_{\text{DMoN}}+W_{\text{dist}}\cdot L_{\text{DPR}}^{\text{Distance}}+W_{\text{var}}\cdot L_{\text{DPR}}^{\text{variance}}+W_{\text{entropy}}\cdot L_{\text{DPR}}^{\text{entropy}}$ (3)
Here, $W_{\text{dist}}$, $W_{\text{entropy}}$, and $W_{\text{var}}$ are hyperparameters that control the influence of each regularization term. The additional terms are designed to promote inter-cluster separation, intra-cluster diversity, and balanced cluster assignments.
@@
where $k$ is the number of clusters, $\mu_{i}$ and $\mu_{j}$ are the centroids of clusters $i$ and $j$, computed as $\mu_{i}=\frac{\sum_{v=1}^{n}C_{vi}X_{v}}{\sum_{v=1}^{n}C_{vi}}$, where $X_{v}$ is the feature vector of node $v$, $\|\cdot\|_{2}$ denotes the Euclidean norm, $\epsilon$ is a predefined threshold that sets the minimum acceptable distance between cluster centroids, and ReLU is the Rectified Linear Unit function, ensuring that only distances below $\epsilon$ contribute to the loss. By penalizing pairs of clusters whose centroids are closer than $\epsilon$, clusters can be effectively pushed apart in the feature space. By encouraging greater separation between clusters, it enhances inter-cluster diversity and reduces overlap, leading to more distinguishable and meaningful clusters.
@@
The variance-based regularization term aims to increase the diversity within each cluster by maximizing the variance of node assignments in each cluster:
@@
[…]
@@
The entropy-based regularization term promotes a uniform distribution of nodes across clusters, preventing any single cluster from dominating. It is formulated as:
@@
where $n$ is the number of nodes, $C_{vi}$ is the soft assignment probability of node $v$ to cluster $i$, $\delta$ is a small constant added to prevent logarithm of zero, ensuring numerical stability. By maximizing the entropy of the cluster assignments, this term encourages uncertainty in the assignments, promoting a more balanced distribution of nodes among clusters. This prevents scenarios where one cluster might contain a majority of the nodes, thus enhancing overall diversity and cluster balance.
Therefore, by integrating these diversity-preserving regularization terms into the DMoN objective, we can simultaneously consider graph structure, cluster balance, inter-cluster separation, and intra-cluster diversity.
@@ 4 Experiments
Datasets
To evaluate the effectiveness of our proposed regularization objective, we conducted experiments on the widely used Cora, CiteSeer, and PubMed datasets Yang et al. (2016), as well as the Coauthor CS and Coauthor Physics datasets (Shchur et al., 2018). The Cora, CiteSeer, and PubMed datasets are citation networks where nodes represent papers and edges denote citation relationships, with node labels corresponding to the topic of each paper. The Coauthor CS and Coauthor Physics datasets are co-authorship graphs, where nodes represent authors and edges indicate collaborations, with labels reflecting the field of research. The key characteristics of these datasets are summarized in Table 1.
Dataset $|V|$ $|E|$ $|X|$ $|Y|$
Cora 2708 5278 1433 7
CiteSeer 3327 4614 3703 6
PubMed 19717 44325 500 3
Coauthor CS 18333 81894 6805 15
Coauthor Phys 34493 247962 8415 5
@@
Following the evaluation protocol outlined in (Tsitsulin et al., 2023), we employed the following metrics to assess performance: graph conductance (C), modularity (Q), Normalized Mutual Information (NMI) with ground-truth labels, and the pairwise F1 measure. These metrics collectively capture the structural quality of clusters and their alignment with ground-truth labels.
Baseline
We select DMoN as our sole baseline because it consistently achieves state-of-the-art results across multiple datasets and metrics, as shown in prior work (Tsitsulin et al., 2023). It outperforms a wide range of established methods—such as k-means(features) (Lloyd, 1982), k-means(DeepWalk) (Perozzi et al., 2014), k-means(DGI) (Veličković et al., 2018), NOCD (Shchur and Günnemann, 2019), DiffPool (Ying et al., 2018), MinCut Pooling (Bianchi et al., 2020), and Ortho (Bianchi et al., 2020)—each representing distinct approaches that prioritize node attributes, graph structure, or both. Unlike traditional clustering methods like k-means (which lacks graph structure awareness) or embedding-based techniques (which emphasize representation learning over explicit clustering), DMoN offers an end-to-end differentiable framework that directly optimizes modularity and uses collapse regularization to avoid trivial solutions. Furthermore, it exhibits superior scalability and computational efficiency compared to pooling-based approaches such as DiffPool and MinCutPool, making it particularly suited for the large, feature-rich datasets in our experiments. By focusing on feature-space diversity within DMoN’s modularity-driven clustering paradigm, we ensure that our enhancements remain aligned with state-of-the-art structural considerations, providing a clear and meaningful evaluation of the improvements introduced by our proposed modifications.
Implementation Details
Similar to (Tsitsulin et al., 2023), one layer of GCN (Kipf and Welling, 2016) with 512 neurons was used to create the graph embedding, followed by the DMoN/DMoN–DPR pooling layer, as illustrated in Figure 1. Likewise, the models were trained for 1000 epochs using Adam optimizer with a learning rate of 0.001. The experiments were run on an Apple M2 Max CPU. The code–implemented by extending the DMoN implementation in PyTorch Geometric (Fey and Lenssen, 2019)– will be made publicly available at www.github.com/TBA.
Figure 1: Overview of the DMoN/DMoN–DPR pipeline.
Hyperparameter Tuning
The weighting coefficients $W_{\text{dist}}$, $W_{\text{entropy}}$, and $W_{\text{var}}$, as well as $\epsilon$ are hyperparameters that need to be tuned based on the specific dataset and desired clustering behavior. They control the trade-off between structural modularity and diversity preservation. To optimize these hyperparameters, we conducted evaluations on the Cora, CiteSeer, PubMed, Coauthor CS and Coauthor Physics datasets using 10 random seeds, similar to how it was done in (Tsitsulin et al., 2023). The seeds that resulted in the best performance when using the DMoN pooling layer were selected. For each dataset, we tuned $\epsilon$ (for the distance weight $W_{\text{dist}}$), $W_{\text{var}}$ (variance weight), and $W_{\text{entropy}}$ (entropy weight) independently. To find the best $\epsilon$, the variance and entropy weights were set to zero, and $\epsilon$ was varied from 10 to $10^{-5}$. Similarly, $W_{\text{var}}$ and $W_{\text{entropy}}$ were tuned by fixing the other weights to zero and varying them across {1, 0.1, 0.01, 0.001}. The best weights identified were used to construct various DMoN-DPR models, including DMoN (baseline), DMoN–DPR(D) (distance), DMoN–DPR(V) (variance), DMoN–DPR(E) (entropy), and combinations: DMoN–DPR(DV) (distance and variance), DMoN–DPR(DE) (distance and entropy), DMoN–DPR(VE) (variance and entropy), and DMoN–DPR(DVE) (distance, variance, and entropy). This systematic approach allowed us to assess the impact of each regularization term and their combinations on model performance. Ablation studies are presented in Section A.

### v2

# Deep Modularity Networks with Diversity-Preserving Regularization

@@ Abstract
Graph clustering plays a crucial role in graph representation learning but often faces challenges in achieving feature-space diversity. While Deep Modularity Networks (DMoN) leverage modularity maximization and collapse regularization to ensure structural separation, they lack explicit mechanisms for feature-space separation, assignment dispersion, and assignment-confidence control. We address this limitation by proposing Deep Modularity Networks with Diversity-Preserving Regularization (DMoN–DPR), which introduces three novel regularization terms: distance-based for inter-cluster separation, variance-based for per-cluster assignment dispersion, and an assignment-entropy penalty with a small positive weight, encouraging more confident assignments gradually. Our method significantly enhances label-based clustering metrics on feature-rich benchmark datasets (paired two-tailed t-test, $p\leq 0.05$), demonstrating the effectiveness of incorporating diversity-preserving regularizations in creating meaningful and interpretable clusters.
@@ 1 Introduction
Graph clustering is a crucial problem within the graph representation learning domain, and essential for various applications, including but not limited to community detection in social networks (Perozzi et al., 2014; Xiao et al., 2015), data exploration (Perozzi and Akoglu, 2018), and functional module identification in biological networks (Jin et al., 2021). Consequently, there has been a surge in methods aimed at enhancing graph clustering performance. Recent advancements have included techniques like Graph Neural Networks (GNNs) (Scarselli et al., 2008), which leverage node features and graph structure for representation learning, often through unsupervised training (Tsitsulin et al., 2023). In this context, graph pooling methods have also become prominent, as they provide a way to coarsen graphs by aggregating nodes into clusters (Cangea et al., 2018). Yet, some of these methods such as DiffPool (Ying et al., 2018) and MinCutPool (Bianchi et al., 2019) were found to be computationally costly and/or too rigid, leading to poor convergence (Tsitsulin et al., 2023).
To address these limitations, Deep Modularity Networks (DMoN) (Tsitsulin et al., 2023) were introduced, combining spectral modularity maximization (Newman, 2006) with collapse regularization (Tsitsulin et al., 2023) to avoid trivial clustering solutions. Although DMoN captures structural communities, two practical gaps remain. First, while the encoder may use node attributes, the objective itself is agnostic to feature‑space geometry: it contains no term that directly rewards inter‑cluster feature separation or specialization of cluster assignments across nodes. This can yield clusters that are structurally distinct yet not well separated in feature space. Second, the objective lacks an explicit control for assignment confidence (e.g., entropy/temperature), so soft assignments can harden early.
Why diversity matters.
Accounting for both inter-cluster separation and intra-cluster variety prevents information collapse and yields richer representations. In practice, diversity improves downstream utility—for example, diversified recommendation in e-commerce optimizes cluster-level variety to boost engagement (Kim and Kang, 2025), and drawing items from distinct clusters raises serendipity without sacrificing accuracy (Aytekin and Karakaya, 2014). Beyond recommendation, biological networks exhibit the same principle: modules that are cohesive inside yet distinct outside reveal functional relationships that density-only objectives can miss (Spirin and Mirny, 2003). Recent graph-pooling work therefore introduces explicit mechanisms to spread centroids and diversify within-cluster features (Liu et al., 2023). These considerations motivate our design.
@@
Motivated by the above, we propose Deep Modularity Networks with Diversity-Preserving Regularization (DMoN-DPR), an extension of DMoN that augments its objective with three regularizers explicitly promoting diversity among clusters. These include a distance-based regularization term, which penalizes clusters with centroids too close in feature space to encourage distinct separation; a variance-based regularization term that increases the dispersion of assignment probabilities across nodes for each cluster, and an assignment‑entropy penalty added with a small positive weight, to avoid premature hard assignments and to let distance/variance terms drive diversity. By performing extensive evaluations, we demonstrate that our method improves clustering performance on the benchmark datasets, noticeably on Coauthor CS and Coauthor Physics datasets, which benefit from this enriched representation due to their rich feature spaces.
@@ 2 Related Work
Early graph clustering often decoupled features from structure. k-means on raw features (Lloyd, 1982) ignores connectivity; pairing k-means with DeepWalk or DGI embeddings (Perozzi et al., 2014; Veličković et al., 2018; Tsitsulin et al., 2023) injects structural signals but remains a two-stage pipeline, typically trailing end-to-end models that jointly learn representations and cluster assignments. Meanwhile, Chebyshev-based spectral convolutions (Defferrard et al., 2016) efficiently approximate graph filters, laying a scalable foundation for modern GNNs.
On the other hand, pooling methods capture hierarchy. NOCD (Shchur and Günnemann, 2019) directly optimizes graph likelihood but can struggle with scale and feature use. DiffPool (Ying et al., 2018) learns soft cluster assignments end-to-end yet incurs quadratic cost (Tsitsulin et al., 2023). MinCut pooling (Bianchi et al., 2020) adds normalized cut and orthogonality terms that may hinder convergence; the Ortho variant keeps only the latter and loses structural cues. SAGPool (Lee et al., 2019) ranks node importance via self-attention, improving selectivity at added training cost on large graphs.
For unsupervised representation learning, DGI (Velickovic et al., 2019) maximizes mutual information between local and global summaries; InfoGraph extends this idea to graph-level representations (Sun et al., 2019).
More recently, DMoN (Tsitsulin et al., 2023) integrates modularity maximization with collapse regularization in an end-to-end framework, unifying community detection and neural feature learning and achieving strong NMI and modularity across benchmarks.
3 Deep Modularity Networks with Diversity-Preserving Regularization
@@
$L_{\text{DMoN}}(C;A)=\underbrace{-\frac{1}{2m}\text{Tr}(C^{\top}BC)}_{\text{Modularity term}}+\underbrace{{\frac{\sqrt{k}}{n}}\left\|\sum_{i}C_{i}^{\top}\right\|_{F}-1}_{\text{Collapse Regularization term}}$ (2)
where $(\frac{1}{2m})\text{Tr}(C^{\top}BC)$ is the modularity term (measuring how well clusters are internally connected compared to random chance), $B=A-\frac{dd^{\top}}{2m}$ is the modularity matrix, $d$ is the degree vector, $m$ is the total number of edges, ${\frac{\sqrt{k}}{n}}\left\|\sum_{i}C_{i}^{\top}\right\|_{F}-1$ represents the collapse regularization term, which discourages the formation of trivial clusters by penalizing the Frobenius norm of the assignment matrix $C$, $\|\cdot\|_{F}$ denotes the Frobenius norm, and $k$ is the number of clusters. This regularization term encourages balanced cluster assignments, thereby improving the quality of clustering and avoiding degenerate solutions.
Building upon the DMoN framework, we augment the objective with three regularizers:
$L_{\text{DMoN-DPR}}(C;A)=L_{\text{DMoN}}+W_{\text{dist}}\,L_{\text{DPR}}^{\text{distance}}+W_{\text{var}}\,L_{\text{DPR}}^{\text{variance}}+W_{\text{entropy}}\,L_{\text{DPR}}^{\text{entropy}}.$ (3)
Here, $W_{\text{dist}}$, $W_{\text{var}}$, and $W_{\text{entropy}}$ control the influence of each term. These additions promote inter-cluster separation (distance) and per-cluster assignment dispersion (variance). We use a small positive entropy weight so per-node entropy decreases slowly, preserving higher uncertainty early and thus indirectly encouraging more balanced cluster usage during exploration; the primary mechanism for cluster-size balance remains DMoN’s collapse regularizer.
@@
where $k$ is the number of clusters, $\mu_{i}$ and $\mu_{j}$ are the centroids of clusters $i$ and $j$, computed as $\mu_{i}=\frac{\sum_{v=1}^{n}C_{vi}X_{v}}{\sum_{v=1}^{n}C_{vi}}$, where $X_{v}$ is the feature vector of node $v$, $\|\cdot\|_{2}^{2}$ denotes the squared Euclidean norm, $\epsilon$ is a predefined threshold that sets the minimum acceptable squared distance between cluster centroids, and ReLU is the Rectified Linear Unit function, ensuring that only distances below $\epsilon$ contribute to the loss. By penalizing pairs of clusters whose centroids are closer than $\epsilon$, clusters can be effectively pushed apart in the feature space. By encouraging greater separation between clusters, it enhances inter-cluster diversity and reduces overlap, leading to more distinguishable and meaningful clusters.
@@
The variance-based regularization term encourages dispersion in the assignment matrix by maximizing, for each cluster, the variance of its assignment probabilities across nodes:
@@
where $\text{Var}(C_{:i})$ denotes the variance of the assignment probabilities of all nodes to cluster $i$. Maximizing this variance prevents uniform cluster columns (e.g., every node assigned with equal probability), ensuring that clusters specialize over different subsets of nodes and thereby enhancing assignment diversity.
@@
We define:
@@
which is the average per‑node Shannon entropy of the soft assignments, where $n$ is the number of nodes, $C_{vi}$ is the soft assignment probability of node $v$ to cluster $i$, $\delta$ is a small constant added to prevent logarithm of zero, ensuring numerical stability. We minimize $L_{\text{DPR}}$ to gently encourage more confident assignments. Because minimizing entropy can lead to premature hardening, we apply a small weight $W_{\text{entropy}}$ (typically 0.001–0.1) so that entropy decreases slowly during training.
@@ 4 Experiments
To evaluate the effectiveness of our proposed regularization objective, we conducted experiments on Cora, CiteSeer, and PubMed datasets (Yang et al., 2016), as well as the Coauthor CS and Coauthor Physics datasets (Shchur et al., 2018). Information regarding these datasets is presented in Appendix A. Following the evaluation protocol outlined in (Tsitsulin et al., 2023), we employed the following metrics to assess performance: graph conductance (C), modularity (Q), Normalized Mutual Information (NMI) with ground-truth labels, and the pairwise F1 measure. Additionally, we use the following quantitative measures to analyze feature‐space diversity: average inter‐centroid distance (mean pairwise Euclidean distance between cluster centroids), minimum inter‐centroid distance (the smallest pairwise distance among centroids), average intra‐cluster variance (average within‐cluster variance of node embeddings), and Silhouette score (a standard measure of per‐point cohesion vs. separation). The quantification of diversity results are presented in Appendix B.
Baseline & Implementation Details
We select DiffPool (Ying et al., 2018), MinCut pooling (Bianchi et al., 2019) and DMoN (Tsitsulin et al., 2023) as our baselines. Additionally, we focus on ablation studies and significance tests against vanilla DMoN to isolate the impact of the new diversity terms. Furthermore, similar to (Tsitsulin et al., 2023), one layer of GCN (Kipf and Welling, 2016) with 512 neurons was used to create the graph embedding, followed by a pooling layer. Likewise, the models were trained for 1000 epochs using Adam optimizer with a learning rate of 0.001. The code–implemented by extending the DMoN implementation in PyTorch Geometric (Fey and Lenssen, 2019)– is available at www.github.com/YasminSalehi/DMoN-DPR.